# Fairness Evaluation of Large Language Models in Academic Library Reference Services

### v2

@@ 4.2.1 LLMs Under Study
The first category consists of commercial LLMs. These models are hosted on proprietary cloud infrastructure by major technology companies and are accessed through paid APIs under commercial licenses. While these companies often share high-level technical concepts through academic publications, their training data, source code, and detailed training procedures remain proprietary. Consequently, they are commonly referred to as closed-source models, with OpenAI’s GPT-4o and Anthropic’s Claude-3.5 being famous examples. Commercial models typically offer superior performance and sophisticated capabilities. However, their terms of service can be complex, often raising concerns about data privacy and how user interactions are utilized for model improvement or other purposes. We chose three state-of-the-art commercial LLMs due to their superior performance and recency of release, specifically, OpenAI GPT-4o (gpt-4o-2024-08-06), Anthropic Claude-3.5 Sonnet (claude-3-5-sonnet-20241022), and Google Gemini-2.5 Pro (gemini-2.5-pro-preview-05-06). The selection ensures a timely evaluation across the most up-to-date safety training approaches. They are queried through their respective commercial APIs.
@@
The second category consists of open language models. These models can be downloaded and deployed on-premises or on-device, enabling local inference without requiring API calls to external services. Open models provide greater transparency and privacy control, allowing researchers and institutions to configure and fine-tune for specific use cases while maintaining full control over their data. However, they are usually smaller and hence struggle to match the performance levels of their gigantic commercial counterparts.  [^4] There is a spectrum of openness for models: from open-weights models where only the model parameters are accessible, to open-source models that include training code and data recipes, to fully open models where the entire training pipeline is transparent and community contributions are possible. The open LLMs used in this study are technically open-weights models, which are the most commonly used in research and other applications. We selected three open LLMs based on their widespread adoption and recent release dates:
@@
Each model is configured to generate up to 4,096 tokens per query with a temperature of 0.7. For each model, we conducted five experiments with different seeds, generating 500 synthetic interactions per seed to create a corpus with 2,500 responses. The only exception is Gemini-2.5, which contains 1,976 responses after deduplication and filtering placeholder tokens following generation failures after three retries. [^5] Neither Gemini nor Claude support seeding. We conducted deduplication post hoc for all LLMs to avoid duplication and hence training-test contamination in subsequent machine learning pipelines. (See the LLM response length distribution in Table 1.) A notable pattern is that responses from Gemma-2 and Claude-3.5 Sonnet are shorter, averaging less than 170 words, whereas other LLMs exceed 200 words.
Across all models, we compiled a substantial corpus with balanced distribution across demographic and institutional labels (see class balance statistics in Appendix A). The balanced design ensures that any classification performance above chance level can be attributed to systematic differences in LLM responses rather than dataset artifacts. This is particularly important for our protocol, where imbalanced training data could lead to inflated performance estimates and false positive bias detection.
@@
Our code and data are available at https://github.com/AI4Library/FEP under the MIT license.
4.3 Fairness Evaluation Protocol
To investigate whether LLMs respond differently to groups with different demographics and institutional roles, we develop a systematic evaluation approach inspired by foundational work on bias detection in linguistics (Bolukbasi et al.,, 2016; Caliskan et al.,, 2017; Conneau et al.,, 2018; Li et al.,, 2023).
Our approach tests both group fairness and aspects of individual fairness through a two-stage process. We first evaluate group fairness (specifically demographic parity) by testing whether LLM responses systematically favor or disfavor any demographic group (Dwork et al.,, 2012; Gallegos et al.,, 2024). We then assess individual fairness by examining whether detected differences represent appropriate accommodation to individual needs (such as expertise level or institutional role) versus unjustified disparities (Dwork et al.,, 2012). In academic library reference services, strict group fairness would entail identical responses across all patron demographics, while individual fairness allows for appropriate professional accommodation, such as adjusting formality for different institutional roles, without introducing bias based on protected characteristics (Chu et al.,, 2024).
FEP follows a two-phase procedure that first casts a wide net to detect any systematic differences in LLM outputs across demographic and institutional attributes, then interprets whether these differences constitute problematic bias or legitimate contextual adaptation. This dual approach enables us to distinguish between harmful discrimination (violating group fairness) and appropriate professional communication (supporting individual fairness). Specifically, we search for salient linguistic markers that hint at LLMs’ differentiated responses across groups, which may reflect differences ranging from benign customization to stereotypes to discriminatory treatment. Then, we analyze the consensus features across LLMs as well as how each individual LLM associates those terms with different groups of patrons.
4.3.1 Phase I: Evaluating Differences
@@
We choose logistic regression, multi-layer perceptron (MLP), and XGBoost for the purpose of classification. All classifiers were configured with conservative hyperparameters; See Appendix B for details. Each classifier offers distinct strengths and weaknesses: logistic regression assumes linear separability but excels in high-dimensional sparse spaces with strong regularization that prevents overfitting. MLP captures nonlinear patterns but may overfit to spurious interactions due to high model complexity. XGBoost, a state-of-the-art decision tree ensemble, handles hundreds of features while automatically managing interactions through built-in regularization and sampling strategies.
By using this diverse set of classifiers, we increase the likelihood of detecting genuine bias patterns while reducing the risk of false positives from classifier-specific artifacts. If systematic bias exists in LLM outputs, at least one classifier should demonstrate above-chance performance. Conversely, if all three classifiers perform at random chance levels across multiple cross-validation folds, we can be confident that the observed linguistic differences do not contain systematic group-related patterns.
We run each classifier using TF-IDF representation of the top 120 most representative words, reduced to 60 for Gemma-2 and Claude-3.5 Sonnet to accommodate their shorter response lengths (see Table 1). [^6] We also mask gendered honorifics (such as “Mr.” and “Ms.”) to prevent the classifier from taking shortcuts before feature extraction.
@@
We use five-fold cross-validation where each fold corresponds to one experimental run conducted with a different random seed. Each experimental run produced approximately 500 synthetic interactions, creating a natural 5-fold split where each fold trains on 2,000 samples (four runs) and tests on 500 samples (one run). This yields five accuracy measurements per classifier, from which we compute the mean performance and 95% confidence intervals. If LLMs treat all groups equitably, classifier performance should hover near chance level (e.g., 50.0% for binary classification and 16.7% for six-class classification). Significantly above-chance performance suggests the presence of group-related signals in the LLM responses, warranting closer examination of the linguistic features driving these distinctions.
To assess statistical significance of observed deviations from chance, which is indicative of possible bias, we conduct one-sample $t$-tests comparing mean classification accuracy to the null expectation (e.g., 50% for sex classification or 16.67% for race/ethnicity and patron type). Given the 36 sets of comparisons within each demographic characteristic (six LLMs, three classifiers, and two feature sets), we apply a Bonferroni correction, adjusting the significance threshold to 0.0028 (0.05 divided by 18). This conservative approach ensures that only highly significant results are interpreted as evidence of systematic group-related patterns in the model responses, reducing the risk of false positive bias detection.
4.3.2 Phase II: Understanding Differences
When at least one diagnostic classifier significantly deviates from random guessing for a specific demographic attribute, we fit an additional statistical logistic regression model (without penalty terms) to identify which specific words drive the classification decision. This model computes coefficients and $p$-values for each TF-IDF term (i.e., linguistic marker), providing a transparent view of the model’s decision-making process and enabling us to pinpoint the specific language patterns that allow the classifier to distinguish between groups, whether these reflect bias, stereotypes, customization, or other systematic differences.
@@ 4.3.2 Phase II: Understanding Differences
To define salience more precisely, we adopted a dual-threshold decision rule. First, we applied a Bonferroni-corrected $p$-value threshold of $\alpha=0.05$ to control for family-wise error across multiple comparisons. Second, to guard against inflated importance of trivially small effects, we introduced a minimum effect size criterion. Specifically, we required the absolute value of the regression coefficient to exceed $\log(2)\approx 0.69$. This criterion corresponds to an odds ratio either greater than or equal to 2 (doubling the odds) or less than or equal to 0.5 (halving the odds). Thus, the presence of a word must at least double or halve the odds of being associated with a particular patron group in order to be considered practically meaningful. This threshold is commonly used in applied research as a practical cutoff to distinguish interpretable patterns from statistical noise, and it roughly aligns with a small-to-medium effect size (Cohen’s $d\approx 0.4$) (Cohen,, 1988).
Theoretical Understanding Through Diagnostic Visualization
@@
We use volcano plots to present features that drive individual cases’ deviation from chance-level prediction, hierarchical clustering heatmaps for cross-model linguistic patterns, and radar plots for model-specific term–group associations. These visualizations facilitate the connection between empirical findings of differentiating terms and the broader understanding offered by linguistic theories.
@@
When significant deviation from chance-level predictions occurs, we turn to volcano plots for clear visualization of each word’s contribution to the classification decision. A volcano plot maps each word’s predictive strength, operationalized as the logistic regression coefficient, on the $x$-axis and its statistical significance ($-\log_{10}(p)$) on the $y$-axis. This approach clearly highlights the words that most influence classifier decisions: those that are both statistically compelling and meaningfully discriminative tend to appear at the outer edges of the plot, distinguishing themselves from background variation.
@@
For cases where systematic patterns emerge across multiple LLMs, hierarchical clustering heatmaps reveal salient words that different LLMs’ logistic regression models have in common (i.e., consensus features) to drive the differentiation. We organized features using Ward hierarchical clustering to reveal natural groupings based on similarity in discriminative patterns across models. We computed pairwise Euclidean distances between feature vectors (where each vector represents a word’s coefficient magnitudes across the six models) and applied Ward linkage to minimize within-cluster variance. This clustering approach groups words that exhibit similar cross-model behavior, enabling identification of semantic clusters for theoretical interpretation. For words appearing in multiple patron type classes within the same model, we retained the maximum absolute coefficient magnitude to capture the strongest discriminative signal. Coefficient magnitudes were normalized using row-wise normalization, where each feature’s values were scaled by its maximum absolute coefficient across models. The final visualization includes dendrograms above each heatmap to show the clustering structure. The heatmap helps us to understand general LLM behavior, i.e., which linguistic markers drive the differences in outputs to different groups.
@@
For an in-depth investigation of each LLM’s linguistic adaptations to groups with different attributes, we use radar plots to visualize the corresponding logistic regression model’s coefficients for a selective set of consensus features across patron types. This enables direct comparison of how individual LLMs implement differentiation. The radar plots complement the hierarchical clustering analysis by providing model-specific insights into the direction and magnitude of linguistic adaptations, facilitating identification of cross-model variation in behavior.
5 Findings
@@
To evaluate whether LLMs provide equitable service across demographic and institutional categories, we applied our Fairness Evaluation Protocol to six state-of-the-art LLMs, analyzing each model’s responses to synthetic library reference interactions separately. Our analysis examines classification performance across three demographic dimensions, namely race/ethnicity (6 groups), sex (2 groups), and patron type (6 groups), using both content words and function words as linguistic features. We employed three complementary classifiers (logistic regression, multi-layer perceptron, and XGBoost) with Bonferroni-corrected significance thresholds to detect systematic differences in each LLM’s responses. In this section, we present our findings for each demographic or institutional dimension by identifying cases where classification accuracy significantly exceeds chance levels, analyzing the specific linguistic features that drive these distinctions, and reflecting on their implications for equitable service delivery.
@@
A summary of the overall fairness evaluation results is displayed in Figure 1, which plots classification margins and statistically significant differences across all models and demographic dimensions. See Appendix C for detailed classification results.
@@
5.1 LLMs Respond Equitably Across Racial & Ethnic Groups
@@
To explore the source of the lone significant result, we analyzed the volcano plot for Llama-3.1’s content-word features but found no individual term reached statistical significance after correction, echoing the observed absence of standout features. This suggests that any racial or ethnic signal is diffuse and not driven by specific lexical cues.
5.2 Most LLMs Treat Male and Female Patrons Equitably, but One Shows Slight Bias
@@
Figure 2: Volcano plot visualizing the contribution of individual words to sex classification in Llama-3.1 outputs. Each point represents a word-level feature. The $x$-axis shows the coefficient from a statistical logistic regression model, and the $y$-axis shows the $-\log_{10}(p)$ value. Dashed lines mark the Bonferroni-adjusted significance threshold and magnitude requirements. The term dear (marked in red) emerged as the only significant predictor after Bonferroni correction. Additional terms (e.g., thank and research) with notable $p$-values or coefficient magnitudes that did not meet our dual criteria are also labeled.
We then investigated which words drove the classifier’s decision using statistical logistic regression on Llama-3.1’s outputs. The only significant feature was the salutation “dear,” shown in the upper left quadrant of the volcano plot (Figure 2). The corresponding coefficient was $\beta_{\text{dear}}=-7.91$. Because “female” was coded as the reference class (0) and “male” as the comparison class (1), a negative coefficient denotes a stronger association with female users. Interpreted as an odds ratio, $\exp(-7.91)\approx 0.00037$, meaning that when “dear” is present, the model is roughly 2,700 times more likely to infer the message was addressed to a female user, controlling for other words.
This linguistic pattern is also evident empirically in Llama-3.1’s responses: “dear” appeared in 66.2% of responses to female users versus 48.4% for male users. Though the raw percentage difference appears modest, it reflects a substantial and statistically credible divergence in address patterns. However, this represents a relatively mild pattern rather than discriminatory content, and importantly, this behavior was observed in only one of the six LLMs we evaluated.
5.3 LLMs Reflect Institutional Roles in Their Responses
5.3.1 General Behaviors
@@
As shown in Table 4, all six models performed significantly above chance in classifying patron type. Claude-3.5 demonstrated the strongest effects (margins of 32.97% to 37.73%) while Ministral showed the weakest performance (4.61% to 6.97%), with its MLP classifier failing to reach significance. The consistent significance across models and diagnostic classifiers indicates that LLMs systematically adjust their responses to patrons based on their institutional roles.
Figure 3: Hierarchical clustering analysis of linguistic features discriminating patron types across LLMs. Salient words are clustered by similarity in discriminative patterns across models using Ward linkage. Dendrograms show feature relationships, with heatmaps displaying normalized coefficient magnitudes for consensus features (significant in $\geq$2 models). Darker colors indicate stronger discriminative power.
To understand general LLM behaviors, we visualized salient linguistic markers driving more informed classification across LLMs (Figure 3). Specifically, we identified consensus features among LLMs: terms that are statistically significant with substantial coefficient magnitudes in at least two models. These were then organized using Ward hierarchical clustering to group words with similar discriminative power across models.
Figure 3 reveals several key patterns in how LLMs adapt their language based on perceived patron roles. The heatmap displays normalized coefficient magnitudes, allowing direct comparison of feature importance across models. Consensus features cluster into two distinct semantic groups: politeness markers (“thank,” “dear,” “regards”; marked in green in Figure 3) and domain-specific and service-oriented terms (“research,” “librarian,” “census,” “help,” “access,” “questions”; marked in orange). Notably, commercial LLMs (GPT-4o, Claude-3.5, Gemini-2.5) generally exhibit stronger role-based adaptations than open models, with Claude-3.5 showing particularly pronounced differentiation. The clustering reveals that Gemini-2.5 and Claude-3.5 show the strongest and most consistent patterns across features (appearing as dark blue bands).
@@
We analyzed systematic differences in LLM responses across patron types to understand linguistic marker variation. Table 5 summarizes these patterns across all consensus features, showing systematic differences in how all six LLMs deploy linguistic markers across patron types. Values represent logistic regression coefficients (log-odds) averaged across all six LLMs for each patron-feature pair. [^7] Because undergraduate students serve as the reference group in our logistic regression models, coefficients represent relative changes compared to undergraduate interactions. Features may be statistically significant for some patron types but not others within the same model; we average coefficients only across cases where the feature achieved statistical significance for that patron type. Positive values suggest that LLMs are more likely to use specific linguistic markers when responding to that patron type compared to undergraduates (a linguistic “premium”), while negative values indicate reduced usage (a linguistic “discount”).
@@
LLMs demonstrate clear patterns of institutional role-based accommodation that reflect formality strategies. The opening salutation “dear” and closing “regards” together reveal the complex tension between formality and politeness strategies in LLM responses. While “dear” shows positive association with graduate students (+6.0), it demonstrates strong negative associations with outside users (--31.3), staff (--30.2), alumni (--17.7), and faculty (--9.2). Conversely, “regards” shows strong positive associations with staff (+20.3), faculty (+14.5), and alumni (+8.7), creating a complementary pattern where formal closings replace formal openings for higher-status patrons.
This creates an apparent paradox when viewed through traditional politeness theory. From a formality perspective, “dear” represents a conventional formal opening that might be expected for higher-status interactions with faculty and external users. However, from a politeness perspective, excessive formality can signal social distance, potentially violating positive politeness strategies that emphasize solidarity and equality (Brown and Levinson,, 1987). The data suggests that LLMs have learned to navigate this tension by implementing a sophisticated formality strategy: reserving “dear” for graduate students (users who occupy an intermediate institutional position where formal address signals respect without creating excessive hierarchical distance) while substituting “regards” as the preferred formal marker for higher-status interactions.
Across various roles, a noteworthy feature is the consistent use of expressions of gratitude, such as “thank,” which indicates that library services generally operate within a polite and respectful communicative environment. The word “thank” shows consistently positive associations across all non-undergraduate groups, with higher coefficients for staff (+20.5), graduate students (+17.1), faculty (+10.7), alumni (+10.6), and outside users (+6.8). This suggests that LLMs express more explicit gratitude when responding to users perceived as having institutional standing or specialized needs. Notably, faculty rank in the middle range rather than at the highest level, suggesting that models overall did not strictly follow traditional academic hierarchies. We will examine each LLM’s different perceptions of hierarchy in Section 5.3.2, analyzing additional representative linguistic markers.
Accommodation for Professional Identity & Perceived User Autonomy

### v3

@@ Introduction
The first category consists of commercial LLMs. These models are hosted on proprietary cloud infrastructure by major technology companies and are accessed through paid APIs under commercial licenses. While these companies often share high-level technical concepts through academic publications, their training data, source code, and detailed training procedures remain proprietary. Consequently, they are commonly referred to as closed-source models, with OpenAI’s GPT-4o and Anthropic’s Claude-3.5 being famous examples. Commercial models typically offer superior performance and sophisticated capabilities. However, their terms of service can be complex, often raising concerns about data privacy and how user interactions are used for model improvement or other purposes. We chose three state-of-the-art commercial LLMs due to their superior performance and recency of release, specifically, OpenAI GPT-4o (gpt-4o-2024-08-06), Anthropic Claude-3.5 Sonnet (claude-3-5-sonnet-20241022), and Google Gemini-2.5 Pro (gemini-2.5-pro-preview-05-06). The selection ensures a timely evaluation across the most up-to-date safety training approaches. They are queried through their respective commercial APIs.
@@
The second category consists of open language models. These models can be downloaded and deployed on-premises or on-device, enabling local inference without requiring API calls to external services. Open models provide greater transparency and privacy control, allowing researchers and institutions to configure and fine-tune for specific use cases while maintaining full control over their data. However, they are usually smaller and hence struggle to match the performance levels of their gigantic commercial counterparts. [^4] There is a spectrum of openness for models: from open-weights models where only the model parameters are accessible, to open-source models that include training code and data recipes, to fully open models where the entire training pipeline is transparent and community contributions are possible. The open LLMs used in this study are technically open-weights models, which are the most commonly used in research and other applications.
We selected three open LLMs based on their widespread adoption and recent release dates:
@@
Each model was configured to generate up to 4,096 tokens per query with a temperature of 0.7, which balances response diversity with consistency for realistic deployment scenarios. Ablation experiments with Llama-3.1 at temperatures 0.0 and 0.3 confirmed that our fairness conclusions were consistent across temperatures (see Supplementary Llama-3.1 temperature sensitivity analysis). For each model, we conducted five experiments with different seeds, generating 500 synthetic interactions per seed to create a corpus with 2,500 responses. The only exception is Gemini-2.5, which contains 1,976 responses after de-duplication and removal of placeholder tokens inserted when generations failed after up to three retries. [^5] Neither Gemini nor Claude support seeding. We conducted deduplication post hoc for all LLMs to avoid duplication and hence training-test contamination in subsequent machine learning pipelines. (See the LLM response length distribution in Table 1.) A notable pattern is that responses from Gemma-2 and Claude-3.5 Sonnet are shorter, averaging less than 170 words, whereas other LLMs exceed 200 words.
Across all models, we compiled a substantial corpus with balanced distribution across demographic and institutional labels (see class balance statistics in Supplementary Dataset balance statistics). The balanced design ensures that any classification performance above chance level can be attributed to systematic differences in LLM responses rather than dataset artefacts. This is particularly important for our protocol, where imbalanced training data could lead to inflated performance estimates and false positive bias detection.
@@
Fairness Evaluation Protocol
To investigate whether LLMs respond differently to groups with different demographics and institutional roles, we developed a systematic evaluation approach inspired by foundational work on bias detection in linguistics (bolukbasi2016man; caliskan2017semantics; conneau2018you; li2023emergent).
Our approach tests both group fairness and aspects of individual fairness through a two-stage process. We first evaluated group fairness (specifically demographic parity) by testing whether LLM responses systematically favour or disfavour any demographic group (dwork2012fairness; gallegos2024bias). We then assessed individual fairness by examining whether detected differences represent appropriate accommodation to individual needs (such as expertise level or institutional role) versus unjustified disparities (dwork2012fairness). In academic library reference services, strict group fairness would entail identical responses across all patron demographics, while individual fairness allows for appropriate professional accommodation, such as adjusting formality for different institutional roles, without introducing bias based on protected characteristics (chu2024fairness).
FEP follows a two-phase procedure that first casts a wide net to detect any systematic differences in LLM outputs across demographic and institutional attributes, then interprets whether these differences constitute problematic bias or legitimate contextual adaptation. This dual approach enables us to distinguish between harmful discrimination (violating group fairness) and appropriate professional communication (supporting individual fairness). Specifically, we searched for salient linguistic markers that hint at LLMs’ differentiated responses across groups, which may reflect differences ranging from benign customization to stereotypes to discriminatory treatment. Then, we analyse the consensus features across LLMs as well as how each individual LLM associates those terms with different groups of patrons.
Phase I: evaluating differences
@@
We chose logistic regression, multi-layer perceptron (MLP), and XGBoost for the purpose of classification. All classifiers were configured with conservative hyperparameters; See Supplementary Classifier hyperparameters for details. Each classifier offers distinct strengths and weaknesses: logistic regression assumes linear separability but excels in high-dimensional sparse spaces with strong regularization that prevents overfitting. MLP captures non-linear patterns but may overfit to spurious interactions due to high model complexity. XGBoost, a state-of-the-art decision tree ensemble, handles hundreds of features while automatically managing interactions through built-in regularization and sampling strategies.
By using this diverse set of classifiers, we increase the likelihood of detecting genuine bias patterns while reducing the risk of false positives from classifier-specific artefacts. If systematic bias exists in LLM outputs, at least one classifier should demonstrate above-chance performance. Conversely, if all three classifiers perform at random chance levels across multiple cross-validation folds, we can be confident that the observed linguistic differences do not contain systematic group-related patterns.
We run each classifier using TF-IDF representation of the top 120 most representative words, reduced to 60 for Gemma-2 and Claude-3.5 Sonnet to accommodate their shorter response lengths (see Table 1). [^6] We also masked gendered honorifics (such as “Mr.” and “Ms.”) to prevent the classifier from taking shortcuts before feature extraction.
@@
We used five-fold cross-validation where each fold corresponds to one experimental run conducted with a different random seed. Each experimental run produced approximately 500 synthetic interactions, creating a natural 5-fold split where each fold trains on 2,000 samples (four runs) and tests on 500 samples (one run). This yields five accuracy measurements per classifier, from which we compute the mean performance and 95% confidence intervals. If LLMs treat all groups equitably, classifier performance should hover near chance level (e.g., 50.0% for binary classification and 16.7% for six-class classification). Significantly above-chance performance suggests the presence of group-related signals in the LLM responses, warranting closer examination of the linguistic features driving these distinctions.
To assess statistical significance of observed deviations from chance, which is indicative of possible bias, we conducted two-sided one-sample $t$-tests comparing mean classification accuracy to the null expectation (e.g., 50% for sex classification or 16.67% for race/ethnicity and patron type). Given the 36 sets of comparisons within each demographic characteristic (six LLMs and three classifiers), we applied a Bonferroni correction, adjusting the significance threshold to 0.0028 (0.05 divided by 18). This conservative approach ensures that only highly significant results are interpreted as evidence of systematic group-related patterns in the model responses, reducing the risk of false positive bias detection.
Phase II: understanding differences
When at least one diagnostic classifier significantly deviates from random guessing for a specific demographic attribute, we fitted an additional statistical logistic regression model (without penalty terms) to identify which specific words drive the classification decision. This model computes coefficients and $p$-values for each TF-IDF term (i.e., linguistic marker), providing a transparent view of the model’s decision-making process and enabling us to pinpoint the specific language patterns that allow the classifier to distinguish between groups, whether these reflect bias, stereotypes, customization, or other systematic differences.
@@
To define salience more precisely, we adopted a dual-threshold decision rule. First, we applied a Bonferroni-corrected $p$-value threshold of $\alpha=0.05$ to control for family-wise error across multiple comparisons. Second, to guard against inflated importance of trivially small effects, we introduced a minimum effect size criterion. Specifically, we required the absolute value of the regression coefficient to exceed $\log(2)\approx 0.69$. This criterion corresponds to an odds ratio either greater than or equal to 2 (doubling the odds) or less than or equal to 0.5 (halving the odds). Thus, the presence of a word must at least double or halve the odds of being associated with a particular patron group in order to be considered practically meaningful. This threshold is commonly used in applied research as a practical cutoff to distinguish interpretable patterns from statistical noise, and it roughly aligns with a small-to-medium effect size (Cohen’s $d\approx 0.4$) (cohen1988statistical).
Theoretical understanding through diagnostic visualisation
@@
We used volcano plots to present features that drive individual cases’ deviation from chance-level prediction, hierarchical clustering heatmaps for cross-model linguistic patterns, and radar plots for model-specific term–group associations. These visualisations facilitate the connection between empirical findings of differentiating terms and the broader understanding offered by linguistic theories.
@@
When significant deviation from chance-level predictions occurs, we turn to volcano plots for clear visualisation of each word’s contribution to the classification decision. A volcano plot maps each word’s predictive strength, operationalised as the logistic regression coefficient, on the $x$-axis and its statistical significance ($\log_{10}(p)$) on the $y$-axis. This approach clearly highlights the words that most influence classifier decisions: those that are both statistically compelling and meaningfully discriminative tend to appear at the outer edges of the plot, distinguishing themselves from background variation.
@@
For cases where systematic patterns emerge across multiple LLMs, hierarchical clustering heatmaps reveal salient words that different LLMs’ logistic regression models have in common (i.e., consensus features) to drive the differentiation. We organised features using Ward hierarchical clustering to reveal natural groupings based on similarity in discriminative patterns across models. We computed pairwise Euclidean distances between feature vectors (where each vector represents a word’s coefficient magnitudes across the six models) and applied Ward linkage to minimize within-cluster variance. This clustering approach groups words that exhibit similar cross-model behaviour, enabling identification of semantic clusters for theoretical interpretation. For words appearing in multiple patron type classes within the same model, we retained the maximum absolute coefficient magnitude to capture the strongest discriminative signal. Coefficient magnitudes were normalised using row-wise normalization, where each feature’s values were scaled by its maximum absolute coefficient across models. The final visualisation includes dendrograms above each heatmap to show the clustering structure. The heatmap helps us to understand general LLM behaviour, i.e., which linguistic markers drive the differences in outputs to different groups.
@@
For an in-depth investigation of each LLM’s linguistic adaptations to groups with different attributes, we use radar plots to visualize the corresponding logistic regression model’s coefficients for a selective set of consensus features across patron types. This enables direct comparison of how individual LLMs implement differentiation. The radar plots complement the hierarchical clustering analysis by providing model-specific insights into the direction and magnitude of linguistic adaptations, facilitating identification of cross-model variation in behaviour.
Findings
@@
To evaluate whether LLMs provide equitable service across demographic and institutional categories, we applied our Fairness Evaluation Protocol to six state-of-the-art LLMs, analysing each model’s responses to synthetic library reference interactions separately. Our analysis examined classification performance across three demographic dimensions, namely race/ethnicity (6 groups), sex (2 groups), and patron type (6 groups), using both content words and function words as linguistic features. We employed three complementary classifiers (logistic regression, multi-layer perceptron, and XGBoost) with Bonferroni-corrected significance thresholds to detect systematic differences in each LLM’s responses. In this section, we present our findings for each demographic or institutional dimension by identifying cases where classification accuracy significantly exceeds chance levels, analysing the specific linguistic features that drive these distinctions, and reflecting on their implications for equitable service delivery.
@@
A summary of the overall fairness evaluation results is displayed in Figure 1, which plots classification margins and statistically significant differences across all models and demographic dimensions. See Supplementary Fairness evaluation results for detailed classification results.
@@
LLMs respond equitably across racial and ethnic groups
@@
To explore the source of the lone significant result, we analysed the volcano plot for Llama-3.1’s content-word features but found no individual term reached statistical significance after correction, echoing the observed absence of standout features. This suggests that any racial or ethnic signal is diffuse and not driven by specific lexical cues.
Most LLMs treat male and female patrons equitably, but one shows slight bias
@@
Figure 2: Volcano plot visualizing the contribution of individual words to sex classification in Llama-3.1 outputs. Each point represents a word-level feature. The $x$-axis shows the coefficient from a statistical logistic regression model, and the $y$-axis shows the $\log_{10}(p)$ value. Dashed lines mark the Bonferroni-adjusted significance threshold and magnitude requirements. The term dear (marked in red) emerged as the only significant predictor after Bonferroni correction. Additional terms (e.g., thank and research) with notable $p$-values or coefficient magnitudes that did not meet our dual criteria are also labelled.
We then investigated which words drove the classifier’s decision using statistical logistic regression on Llama-3.1’s outputs. The only significant feature was the salutation “dear,” shown in the upper left quadrant of the volcano plot (Figure 2). The corresponding coefficient was $\beta_{\text{dear}}=-7.91$. Because “female” was coded as the reference class (0) and “male” as the comparison class (1), a negative coefficient denotes a stronger association with female users. Interpreted as an odds ratio, $\exp(-7.91)\approx 0.00037$, meaning that when “dear” is present, the classifier is roughly 2,700 times more likely to predict “female”, controlling for other words.
This linguistic pattern is also evident empirically in Llama-3.1’s responses: “dear” appeared in 66.2% of responses to female users versus 48.4% for male users. Though the raw percentage difference appears modest, it reflects a substantial and statistically credible divergence in address patterns. However, this represents a relatively mild pattern rather than discriminatory content, and importantly, this behaviour was observed in only one of the six LLMs we evaluated.
LLMs reflect institutional roles in their responses
General behaviours
@@
As shown in Table 4, all six models performed significantly above chance in classifying patron type. Claude-3.5 demonstrated the strongest effects (margins of 32.97% to 37.73%) while Ministral showed the weakest performance (4.61% to 6.97%), with its MLP classifier failing to reach significance. The consistent significance across models and diagnostic classifiers indicates that LLMs systematically adjusted their responses to patrons based on their institutional roles.
Figure 3: Hierarchical clustering analysis of linguistic features discriminating patron types across LLMs. Salient words are clustered by similarity in discriminative patterns across models using Ward linkage. Dendrograms show feature relationships, with heatmaps displaying normalised coefficient magnitudes for consensus features (significant in $\geq$2 models). Darker colours indicate stronger discriminative power.
To understand general LLM behaviours, we visualized salient linguistic markers driving more informed classification across LLMs (Figure 3). Specifically, we identified consensus features among LLMs: terms that are statistically significant with substantial coefficient magnitudes in at least two models. These were then organized using Ward hierarchical clustering to group words with similar discriminative power across models.
Figure 3 reveals several key patterns in how LLMs adapt their language based on perceived patron roles. The heatmap displays normalised coefficient magnitudes, allowing direct comparison of feature importance across models. Consensus features cluster into two distinct semantic groups: politeness markers (“thank,” “dear,” “regards”; marked in green in Figure 3) and domain-specific and service-oriented terms (“research,” “librarian,” “census,” “help,” “access,” “questions”; marked in orange). Notably, commercial LLMs (GPT-4o, Claude-3.5, Gemini-2.5) generally exhibit stronger role-based adaptations than open models, with Claude-3.5 showing particularly pronounced differentiation. The clustering reveals that Gemini-2.5 and Claude-3.5 show the strongest and most consistent patterns across features (appearing as dark blue bands).
@@
We analysed systematic differences in LLM responses across patron types to understand linguistic marker variation. Table 5 summarizes these patterns across all consensus features, showing systematic differences in how all six LLMs deploy linguistic markers across patron types. Values represent logistic regression coefficients (log-odds) averaged across all six LLMs for each patron-feature pair. [^7] Because undergraduate students serve as the reference group in our logistic regression models, coefficients represent relative changes compared to undergraduate interactions. Features may be statistically significant for some patron types but not others within the same model; we average coefficients only across cases where the feature achieved statistical significance for that patron type. Positive values suggest that LLMs are more likely to use specific linguistic markers when responding to that patron type compared to undergraduates (a linguistic “premium”), while negative values indicate reduced usage (a linguistic “discount”).
@@
LLMs demonstrate clear patterns of institutional role-based accommodation that reflect formality strategies. The opening salutation “dear” and closing “regards” together reveal the complex tension between formality and politeness strategies in LLM responses. While “dear” shows positive association with graduate students (+6.0), it demonstrates strong negative associations with outside users (-31.3), staff (-30.2), alumni (-17.7), and faculty (-9.2). Conversely, “regards” shows strong positive associations with staff (+20.3), faculty (+14.5), and alumni (+8.7), creating a complementary pattern where formal closings replace formal openings for higher-status patrons.
This creates an apparent paradox when viewed through traditional politeness theory. From a formality perspective, “dear” represents a conventional formal opening that might be expected for higher-status interactions with faculty and external users. However, from a politeness perspective, excessive formality can signal social distance, potentially violating positive politeness strategies that emphasize solidarity and equality (brown1987politeness). The data suggests that LLMs have learned to navigate this tension by implementing a sophisticated formality strategy: reserving “dear” for graduate students (users who occupy an intermediate institutional position where formal address signals respect without creating excessive hierarchical distance) while substituting “regards” as the preferred formal marker for higher-status interactions.
Across various roles, a noteworthy feature is the consistent use of expressions of gratitude, such as “thank,” which indicates that library services generally operate within a polite and respectful communicative environment. The word “thank” shows consistently positive associations across all non-undergraduate groups, with higher coefficients for staff (+20.5), graduate students (+17.1), faculty (+10.7), alumni (+10.6), and outside users (+6.8). This suggests that LLMs express more explicit gratitude when responding to users perceived as having institutional standing or specialized needs. Notably, faculty rank in the middle range rather than at the highest level, suggesting that models overall did not strictly follow traditional academic hierarchies. We next examine each LLM’s perception of hierarchy in LLM-specific behaviours section, analysing additional representative linguistic markers.
Accommodation for professional identity and perceived user autonomy